Use variation of parameters:

4. (D^2 + 2D +2)y = (e^-x) csc(x)
Answer #1

112 9 The given differential equation is (B? +2D+afy = ē y coseclor) - 1 Where D = doc Egyceffor (iis a linear differential e.: The complimeutty function (CoFJ PJ of equation (1) is given by CoF = Ora la Cos B4+ą SMBA) * = ēl ci conat Disa), 6:2=-1,B13) functious and w=wrous kian of the - S, and yo na w=1 - yi you W= le coura) ēt since 1 ē99-ecospa Cosa-ētsino =ē Cestayf ēS 2 A = 5 (-es) gulay cosecca) de les =-347, 310 Coseco = 13 so we have A = -4 - 14 Again from 17), (11), (1.2) and (13), we1900 Hence we have Yb= ē% - 0 Cosa) + sin(a) clog S4 (213—116) Is the particular sodation of the given differential equation